If the trip home takes 39 minutes, at what time does Chris have to leave in order to be home by a quarter after 6?
Found 2 solutions by stanbon, josmiceli:
Answer by stanbon(75887)

If the trip home takes 39 minutes, at what time does Chris have to leave in order to be home by a quarter after 6?
---------------------------
Home by 6:15
----
6:15 - 39 minutes = 6:00 - 24 minutes = 5:36
-------
Answer: Chris should leave by 5:36 to get home by 6:15
=======================
Cheers,
Stan H.
=======================

Answer by josmiceli(19441)

Chris has to leave 39 minutes before 6:15
1 hr 15 min = 75 min
75 - 39 = 36
Chris has to leave at 5:36
-----------
check:
6:00 - 5:36 = 24 min
24 + 15 = 39 min for the trip